OpenStudy (anonymous):

Which Lemonade mix has a stronger taste? a.) 4 cups of water- 8/3 spoons of lemon b.) 2 cup of water - 1/3 spoons of lemon c.) 10 cups of water - 8 spoons of lemon d.) 30 cups of water - 20 spoons of lemon I know the answer is c but how did they get the answer? Ps. I have problems with my internet so if I disconnect I'll be trying to connect again as fast as I can. Sorry if I make anyone wait.

OpenStudy (anonymous):

see 10/8=1.25 less water,4/8/3=1.5,30/20=1.5,2/1/3=6

OpenStudy (anonymous):

r u there

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

so the one that equals less water its the strongest taste

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh ok it makes sense

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

yw